OpenStudy (anonymous):

The Kw of water: []varies with ph []is only applicable to neautral water []remains constant regardless of the ph [] ranges between 1.0 and 10-14

OpenStudy (anonymous):

It has nothing to do with pH, so the third choice is correct.

OpenStudy (anonymous):

\[K_{eq} = \frac{[H_3O^+] [OH^-]}{[H_2O]^2} \] \[K_W = [H_3O^+][OH^-] = K_{eq} \times [H_2O]^2 = 1 \times 10^{-14} @ 25^oC\]

OpenStudy (anonymous):

The does indeed make choice C correct, although it DOES change depending on temperature. For example, dissolving a strong acid or base into water will change the temperature of the water as it dissociates, which is why you do it slowly (and never put the water directly into concentrated strong acids or bases).

OpenStudy (anonymous):

The pH measures the concentration of [H\(_3\)O\(^+\)]. Do you see why now changing it around means the constant Kw doesn't change? It's like a seesaw, with Kw being the fulcrum at the center.

OpenStudy (anonymous):

yeah that's right